OpenStudy (anonymous):

Checking my work :) 2z^-4 I have to rewrite without the negative exponent. So I got 1/2z^4 is this correct?

OpenStudy (sweetburger):

well if the 2z were in parenthesis likes this (2z)^-4 then you would go ahead and flip the 2 aswell but the in this case the ^-4 only pertains to the variable z

OpenStudy (sweetburger):

how do you think your answer should change?

OpenStudy (sweetburger):

any questions @SeasAventura

OpenStudy (anonymous):

so I'm not supposed to keep my two at the bottom like I would usually do with the negative exponents formula?

OpenStudy (shamim):

u know thay x^-1=1/x

OpenStudy (shamim):

bt 2x^-1=2/x

OpenStudy (anonymous):

Ah so I multiply the two against the one and that's my numerator. Gotcha.

OpenStudy (shamim):

(2x)^-1=1/(2x)

OpenStudy (anonymous):

Thank you

OpenStudy (shamim):

wc

OpenStudy (shamim):

who will get medal?

OpenStudy (sweetburger):

its really coming down to the medals lol

OpenStudy (sweetburger):

lol it doesnt matter